OpenStudy (anonymous):

how would you solve this? 7/radical 12 ? i know you multiply both sides by radical 12 but i dont know what to do after this?

hartnn (hartnn):

both sides ? you mean numerator and denominator ?

hartnn (hartnn):

so what you have in denominator ?

OpenStudy (anonymous):

the denominator is radical 12 and yeah thats what i meant

hartnn (hartnn):

after you multiply num and denom by sqrt 12 you will have \(\sqrt {12}\times \sqrt {12}\) in the denominator, right ?

hartnn (hartnn):

correct! simplify denominator and you're done hint : \(\sqrt x \times \sqrt x =x\)

OpenStudy (anonymous):

would the bottom just equal to 12 now?

hartnn (hartnn):

yes!

OpenStudy (anonymous):

would \[\sqrt{12} = 3\sqrt{2}\sqrt{2}\]

hartnn (hartnn):

no.... where did the square root for 3 go ?

hartnn (hartnn):

\(\sqrt {12}= \sqrt {3\times 4} = \sqrt 3 \sqrt 4 =... ?\)

hartnn (hartnn):

and whats the square root of 4 ?

OpenStudy (anonymous):

2

hartnn (hartnn):

so, \(\sqrt {12}=2\sqrt3\)

hartnn (hartnn):

so everything clear till here ? \(\dfrac{7\times 2\sqrt3}{12}\) ?

hartnn (hartnn):

what gets cancelled out ?

hartnn (hartnn):

you still there ? or did i scare you off ?